### v2

# Natural super-orbitals representation of many-body operators

Abstract
We introduce the concept of natural super-orbitals for many-body operators, defined as the eigenvectors of the one-body super-density matrix associated with a vectorized operator. We relate these objects to measures of non-Gaussianity of operators associated to the occupations of the natural super-orbitals, and define how the non-stabilizerness of operators can be affected by such a basis rotation. We first analyze the general analytical properties of these objects in various contexts, including the time-evolution operator of non-interacting Hamiltonians and Haar-random unitaries. We then perform a numerical investigation of the natural super-orbitals corresponding to both the time-evolution operator and a time-evolved local operator, focusing on two many-body systems: the fermionic $t\text{-}V$ chain and an impurity model, using tensor network simulations. Our results reveal that the $t\text{-}V$ model lacks a preferred super-orbital basis, while in the impurity model, the occupations of the natural orbitals for both operators decay exponentially at all times. This indicates that only a small number of orbitals contribute significantly to quantum correlations, enabling a compact matrix-product-operator representation and a reduced non-stabilizerness in the natural orbital basis. Finally, we examine the spatial spread of the natural orbitals for time-evolved local operators in the impurity model and show that the complexity of this operator in the natural orbital basis saturates over time. This new framework opens the door to future research that leverages the compressed structure of operators in their natural super-orbital basis, enabling for instance the computation of out-of-time-order correlators in large interacting systems over extended time scales.
I Introduction
Strong correlations between particles, arising from their interactions, are at the heart of a wide array of exotic phenomena in condensed matter physics, ranging from high-temperature superconductivity [1] to quantum spin liquids [2] and many-body localization [3, 4]. Understanding the physics of such strongly correlated systems is notoriously difficult, as it typically requires handling exponentially large Hilbert spaces—especially in non-perturbative regimes where analytic approaches fail. To overcome these challenges, a wide range of numerical techniques have been developed over the past decades, including exact diagonalization [5, 6], quantum Monte Carlo [7, 8], and tensor network algorithms [9, 10]. Among these, tensor network methods—particularly due to the efficient and practical matrix product states (MPS) and matrix product operators (MPO) representations—have become a method of choice for tackling one-dimensional systems where entanglement remains sufficiently bounded.
Operators play a central role in all these approaches. They encompass observables used to extract physical quantities, time-evolution operators that govern the dynamics of states or operators (in either the Schrödinger or Heisenberg picture), and the Hamiltonian itself, which is the cornerstone of simulations such as density matrix renormalization group (DMRG) [11, 12, 9]. Moreover, operators can encode information about a quantum system that is inaccessible from the states alone—for instance, through out-of-time-order correlators (OTOCs), which provide insight into the scrambling of quantum information across space and time [13, 14, 15, 16, 17, 18].
In tensor network simulations, operators are typically encoded as MPOs. The efficiency of such representations depends critically on the internal correlations of the operator: higher correlations translate into larger bond dimensions and computational overhead. This is especially true for unitary operators, such as the time-evolution operator or time-evolved local operators, whose von Neumann entanglement entropy tends to grow rapidly—following either a volume or area law—thus making their simulation particularly challenging [19, 20, 21, 22, 23, 18]. This growth is even more severe than in the MPS case, as the operator acts on a doubled Hilbert space, effectively squaring the system size and amplifying entanglement barriers.
A crucial but often under-explored aspect of MPS and MPO simulations is the choice of the one-particle basis. The complexity of these tensor network representations is not intrinsic to the physical system alone, but also depends on how the degrees of freedom are organized. Specifically, it is desirable to identify an orbital basis in which the correlations—and hence the entanglement—between subsystems are minimized. Recent advances have demonstrated that, for specific problems such as quantum impurity models, the entanglement growth during a quench can be overcome by implementing a suitable time-dependent basis rotation [24]. In these systems, consisting of a strongly interacting impurity coupled to a large non-interacting fermionic bath, long-range correlations emerge naturally. Yet, it has been shown that expressing states in terms of natural orbitals—eigenvectors of the one-body density matrix—enables significant compression, due to the exponential decay of their occupation in both ground states [25, 26, 27, 28, 29, 30, 31] and certain out-of-equilibrium settings [24]. Natural orbitals have also proven useful in characterizing many-body localized systems, where different basis choices reveal distinct one-particle properties [32, 33, 34, 35].
Building on this idea, we explore whether similar reductions in complexity can be achieved directly at the level of operators by introducing the concept of natural super-orbitals associated to operators. We define these as the eigenvectors obtained from a vectorized form of an operator and its corresponding one-body super-density matrix, a natural generalization of the concept used for quantum states. Inspired by the structure of many-body states in particular physical contexts, we investigate the properties of unitary operators in spin chains and quantum impurity problems, expecting that the behavior of their natural super-orbitals will vary significantly between these two settings. Our main result shows that, in quantum impurity models, the complexity of MPO representations can often be exponentially reduced by working in the natural super-orbital basis of the operator. This suggests that the compression seen in specific states (e.g., ground states) actually stems from fundamental properties of operator themselves—properties that extend across the full many-body Hilbert space.
Beyond its utility in quantifying and potentially reducing the complexity of classical simulations, the framework presented here also proves valuable for characterizing computational complexity in quantum simulations. A central topic of recent research has been the classification and quantification of quantum operations based on their resource requirements [36]. According to the Gottesman-Knill theorem, Clifford circuits can be simulated efficiently on classical computers, as they require only polynomial resources [37, 38]. However, these circuits cannot generate all quantum states. To achieve universal quantum computation, non-Clifford gates must be introduced, which significantly increases the difficulty of classical simulations [39, 40]. The complexity introduced by such non-Clifford operations is captured by the concept of non-stabilizerness [41, 42, 43, 44], also known as quantum magic, which has also can also be extended to fermionic systems [45, 46]. Quantifying this resource demands efficient and practical methods for classical computation [47, 48, 49, 50, 51]. Similarly, fermionic Gaussian states—eigenstates of non-interacting Hamiltonians or those generated by matchgate circuits—are also amenable to efficient classical simulation [52, 53]. In contrast, the inclusion of interactions introduces non-Gaussian features, thereby increasing complexity as the underlying computational Fock space is exponentially large. This other kind of complexity is captured by the measure of non-Gaussianity, which is interesting for different class of problems [54, 55, 56, 57, 46]. Both non-stabilizerness and non-Gaussianity become particularly meaningful in the context of operators, which represent quantum circuits or time-evolution operators associated to a given Hamiltonian [53, 58], and have been so far less investigated due to the lack of framework for operators. In this work, we demonstrate that these complexity measures can be extended to many-body operators and efficiently computed using tensor network methods. We also explore the properties of these measures in the specific systems studied here.
The paper is organized as follows: In Section II, we review the properties of the one-body density matrix for quantum states, focusing on the features we aim to generalize to operators. In Section III, we introduce the notion of natural super-orbitals for operators and define orbital rotations in terms of super-orbitals in a doubled Hilbert space. We also analyze the properties of such rotations when the operator is unitary, and illustrate the formalism with a simple case of a factorized operator. We then discuss the measures of non-Gaussianity and non-stabilizerness for operators. In Section IV, we consider two limiting cases: non-interacting systems, where the natural orbital basis trivializes the MPO structure, and Haar-random unitaries, where no basis is preferred. The main numerical results are presented in Section V, where we investigate the time-evolution operator and time-evolved local observables in the ttVV model and in a quantum impurity model. We show that in the latter case, the MPO complexity as well as the basis-dependent non-stabilizerness can be significantly reduced using the natural orbital basis. Finally, in Section VI, we summarize our findings and outline possible future directions stemming from the concepts introduced in this work.
II One-body density matrix and natural orbitals for states
We start this article by reviewing the definition of the usual one-body density matrix defined for an arbitrary state $|\psi\rangle$, and some of its properties that we extend for operators later on. Given a fermionic system in which the particle can occupy $L$ orbitals of any kind (we only consider orbitals with occupation zero or one here), equipped with the fermionic creation and annihilation operators $\hat{c}^{\dagger}_{l}$ and $\hat{c}_{l}$ respectively creating and removing a fermion in the orbital $l$, a state $|\psi\rangle$ can be written in a complete basis of the fermionic Fock space:
$|\psi\rangle=\sum_{i=1}^{2^{L}}a_{i}|i\rangle=\sum_{i=1}^{2^{L}}a_{i}\left[\prod_{l=1}^{L}\left(\hat{c}^{\dagger}_{l}\right)^{n^{i}_{l}}\right]|0\rangle,$ (1)
where $a_{i}$ are complex coefficients which for a normalized state are constrained as $\sum_{i}|a_{i}|^{2}=1$ and $|i\rangle$ are product states forming a complete basis $\sum_{i}|i\rangle\langle i|=\mathds{1}$. $n^{i}_{l}=\langle i|\hat{n}_{l}|i\rangle$ is the occupation of the orbital $l$ in the state $|i\rangle$ and $|0\rangle$ is the vacuum state. The one-body density matrix, which we label here $Q$, is a $(L\times L)$ hermitian matrix (and not an operator like the state density matrix), whose elements are defined for such a state as follows:
$Q_{ij}=\langle\psi|\hat{c}^{\dagger}_{i}\hat{c}_{j}|\psi\rangle.$ (2)
$Q$ being hermitian, there exist a unitary matrix $V$ such that $Q=VDV^{\dagger}$, such that $V$ corresponds to a rotation of the one-particle basis of orbitals $\hat{c}$ in the so-called natural orbitals basis $\hat{q}$. The basis rotation can be written as a linear combination of the first set of orbitals with the elements of $V$:
$\hat{q}_{\alpha}=\sum_{i=1}^{L}V^{*}_{i\alpha}\hat{c}_{i},\quad\quad\hat{c}_{i}=\sum_{\alpha=1}^{L}V_{i\alpha}\hat{q}_{\alpha}.$ (3)
The elements of the diagonal matrix $D$ are the eigenvalues of $Q$, which are the occupations of the natural orbitals $\hat{q}$ in the state $|\psi\rangle$, namely:
$D_{\alpha,\beta}=\langle\psi|\hat{q}^{\dagger}_{\alpha}\hat{q}_{\beta}|\psi\rangle\delta_{\alpha,\beta}=n_{\alpha}\delta_{\alpha,\beta}.$ (4)
The property of the natural orbitals is that it is the basis in which the sum of the occupations $\sum_{a=1}^{r}n_{a}$ of the first $r$ orbitals sorted in descending order is maximized. This object is widely used in the quantum chemistry community to organize molecular orbitals according to their degree of correlations [59, 60, 61]. Therefore, if the state $|\psi\rangle$ can be represented as a product state in a given (known or unknown) one-particle basis, it will be revealed by the natural orbital basis, whose occupations will be strictly $n_{\alpha}=1$ or 0, if the orbital $\alpha$ is occupied by a fermion in the state ($n_{\alpha}=1$) or not ($n_{\alpha}=0$). Correlations lead to occupations that are different from zero or one.
III Natural super-orbitals for operators
The matrix $Q$ that we defined above serves as a tool to assess the complexity of a given state by measuring its proximity to a product state. We now extend this line of inquiry to operators: can we construct an analogous object that captures the complexity of an operator—specifically, one that indicates whether the operator can be expressed as a product operator, with a suitable definition of what constitutes a product operator? This question is particularly relevant for the numerical simulation of quantum systems in the Heisenberg picture, as discussed in the introduction. In this framework, evolving operators are often represented as matrix product operators (MPOs), whose complexity depends on the chosen basis. A notable example is that of free systems, where the time-evolution operator can be expressed as an MPO with bond dimension one in the orbitals basis in which the Hamiltonian matrix is diagonal, but it becomes unmanageable in any other one-particle basis.
In order to follow the same derivation as before, it is always possible to vectorize operators such that their mathematical formulation is similar to the one for states. The vectorization of an operator $\hat{O}$ is usually defined through the following isomorphism:
$\begin{split}\hat{O}&=\sum_{i,j=1}^{2^{L}}O_{ij}|i\rangle\langle j|,\\ |\hat{O}\rangle\!\rangle&:=\sum_{i,j=1}^{2^{L}}O_{ij}|i\rangle|j\rangle.\end{split}$ (5)
Following the form of a state as defined in Eq. (1), we can write an arbitrary operator $\hat{O}$ using creation and annihilation operators:
$\begin{split}\hat{O}&=\sum_{i,j=1}^{2^{L}}O_{ij}\left[\prod_{l=1}^{L}\left(\hat{c}^{\dagger}_{l}\right)^{n^{i}_{l}}\right]|0\rangle\langle 0|\left[\prod_{l=L}^{1}\left(\hat{c}_{l}\right)^{n^{j}_{l}}\right],\\ |\hat{O}\rangle\!\rangle&=\sum_{i,j=1}^{2^{L}}O_{ij}\left[\prod_{l=1}^{L}\left(\hat{c}^{\dagger}_{l}\right)^{n^{i}_{l}}\right]\otimes\left[\prod_{l=1}^{L}\left(\hat{\tilde{c}}_{l}\right)^{n^{j}_{l}}\right]^{T}\!\!\!|0\rangle\!\rangle,\end{split}$ (6)
where the transpose is taken on the operators acting on the dual space labelled with a tilda, since they act from the left to the right in Eq. (5). The vacuum state is the tensor product of the vacua of the direct and dual spaces, $|0\rangle\!\rangle=|0\rangle\otimes|0\rangle$, or equivalently the vacuum of a super-space corresponding to a doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$. We can simplify the form of Eq. (6), by introducing the following super-operators acting on the doubled Hilbert space, which we label with a double hat:
$\hat{\hat{d}}^{\dagger}_{m}=\begin{cases}\hat{c}^{\dagger}_{m}\otimes\mathds{1}\hskip 41.25641pt\text{if }\quad m\leq L\\ \mathds{1}\otimes\hat{\tilde{c}}_{m-L}\hskip 28.45274pt\text{otherwise}\end{cases},$ (7)
where $m\in[1,2L]$ and $\hat{\tilde{c}}$ means that the matrix representation of the operator is transposed, since it should be applied from its left and not from its right as usual. Using these super-operators, the vectorized operator from Eq. (6) can be written:
$|\hat{O}\rangle\!\rangle=\sum_{k=1}^{4^{L}}O_{k}\left[\prod^{2L}_{m=1}\left(\hat{\hat{d}}_{m}^{\dagger}\right)^{n^{k}_{m}}\right]|0\rangle\!\rangle,$ (8)
which is now very similar to the state in Eq. (1). The former indices $i$ and $j$ from Eq. (6) can be recovered from the reshape of the matrix $O_{ij}$ of shape $(2^{L}\times 2^{L})$ into the vector $O_{k}$ of dimension $4^{L}$. A graphical notation for the vectorization of an operator can be the following:
where $h_{i}$ (resp. $\tilde{h}_{i}$) refers to the local direct (resp. dual) Hilbert space at site $i$, associated with the gray (resp. orange) color.
The introduction of the super-operators $\hat{\hat{d}}_{m}$ allows us to extend the definition of the one-body density matrix to operators in terms of their vectorized form and the $2L$ super-operators of Eq. (7). We call $R[\hat{O}$] the $(2L\times 2L)$ dimensional hermitian super-density matrix associated to the vectorized form of the operator $\hat{O}$, whose elements read:
[…]
which naturally takes the form of a matrix with four blocks, the off-diagonal onces mixing the direct and dual spaces of the operator. As for the matrix $Q$, $R$ can be diagonalized by a unitary matrix $W$ such that $R=WDW^{\dagger}$, which also represents a rotation of the super-orbitals, forming the basis of super-natural-orbitals:
$\hat{\hat{q}}_{\alpha}=\sum_{m=1}^{2L}W^{*}_{m\alpha}\hat{\hat{d}}_{m},\quad\quad\hat{\hat{d}}_{m}=\sum_{\alpha=1}^{2L}W_{m\alpha}\hat{\hat{q}}_{\alpha}.$ (10)
The specificity of the rotation $W$ compared to the previous rotation $V$ comes from the off-diagonal blocks of $R$, and is not simply a rotation of the system’s orbitals $\hat{c}$ but also mixes the dual and direct spaces of the operator. The occupation $n_{\alpha}=\langle\!\langle\hat{O}|\hat{\hat{q}}_{\alpha}^{\dagger}\hat{\hat{q}}_{\alpha}|\hat{O}\rangle\!\rangle$ of the super-orbitals in the vectorized operator are also given by the elements of the diagonal matrix $D$, corresponding to the eigenvalues of $R$. In the following, we refer to these eigenvectors as the correlation spectrum of the operator. Using the graphical notation for tensors, the elements of the matrix $R$ can be computed as depicted in the following sketch:
The one-body super-density matrix of an operator as defined in Eq. (9) has been already introduced in Refs. [23, 62, 63] as a conformal field theory tool to study bipartite entanglement properties, but no attention has been put to these objects as measures of the complexity of operators. As stated for the case of states, the natural super-orbitals of the operator can reveal if there exist (or not) a basis in which the operator can be expressed as a product operator, for which Eq. (8) has only a single element. However, due to the mixing of the direct and dual spaces, the operator might not be an MPO anymore but an MPS in the doubled Hilbert space, unless the natural super-orbitals preserve the tensor product structure between the direct and the dual space. In practice, the matrix $R$ can be computed efficiently without having to vectorize the operator, by noticing the following equivalences:
[…]
such that elements of $R$ can be efficiently computed by computing the trace of the product of MPOs, which is a common operation in tensor network simulations. Note that for states, the normalization condition is such that $\langle\psi|\psi\rangle=1$, which is not the case for operators anymore. Thus, we normalize the elements of $R$ by the norm of the operator, which reads $\langle\!\langle\hat{O}|\hat{O}\rangle\!\rangle=\mathrm{Tr}[\hat{O}^{\dagger}\hat{O}$].
In the following, we present some properties of the matrix $R$ for specific types of operators, namely unitary operators and density matrices, and we show a simple example of the meaning of a product operator in the sense presented above.
Unitary operators
Unitary operators have the property that $\hat{O}^{\dagger}\hat{O}=\mathds{1}$, which leads to some simplifications in the definition of the operator-one-body density matrix as is clear from Eqs. (11). By using the cyclicity of the trace, we can rewrite the following elements as:
[…]
where trace of the identity disappears with the division by the norm of the operator. Thus, only off-diagonal blocks of the matrix $R$ are non-trivial and need to be computed for unitary operators. The matrix $R$ has then the following form:
$R=\begin{pmatrix}\frac{1}{2}\mathds{1}_{L}&A\\ A^{\dagger}&\frac{1}{2}\mathds{1}_{L}\\ \end{pmatrix}=\frac{\mathds{1}_{2L}}{2}+\begin{pmatrix}0&A\\ A^{\dagger}&0\\ \end{pmatrix}$, (13)
where $A_{mn}=\mathrm{Tr}(\hat{O}^{\dagger}\hat{c}^{\dagger}_{m}\hat{O}\hat{c}_{n})$ is the only-non trivial $(L\times L)$ block of $R$. With such a form, the eigenvalues of $R$ can be found through the eigenvalues of $A^{\dagger}A$ only. If we denote the eigenvalues of $A^{\dagger}A$ as $\mu_{\alpha}$, the eigenvalues of $R$ come by pairs as $\lambda_{\alpha}^{\pm}=\frac{1}{2}\pm\sqrt{A^{\dagger}A}=\frac{1}{2}\pm|\mu_{\alpha}|$. If $A$ is normal (hermitian, unitary, etc. such that $A^{\dagger}A=AA^{\dagger}$), then $\mu_{\alpha}$ are the eigenvalues of $A$. The eigenvalues of $R$ being comprised between 0 and $1$, the eigenvalues of $A$ are thus restricted to the interval $\frac{1}{2},\frac{1}{2}]$. The eigenvectors of $R$ are just two copies of the eigenvectors of $A^{\dagger}A$ with a sign depending on the corresponding eigenvalue one for each off-diagonal block of the matrix: $W_{\alpha,\pm}=\frac{1}{\sqrt{2}}[v_{\alpha},\pm v_{\alpha}]$, where $v_{\alpha}$ is the eigenvector of $A^{\dagger}A$ associated to the eigenvalue $\mu_{\alpha}$. We recall that the doubled Hilbert space is divided into the operators acting on the direct space first, then on the dual space, as $\{\hat{c}^{\dagger}_{1},\hat{c}^{\dagger}_{2},\cdots,\hat{\tilde{c}}_{1},\hat{\tilde{c}}_{2},\cdots\}$. With the eigenvalues coming in pairs, the eigenvectors also come in pairs:
$\hat{\hat{q}}^{\dagger}_{\alpha,\pm}=\sum_{i=1}^{L}\frac{v_{i\alpha}^{*}}{\sqrt{2}}\left(\hat{c}^{\dagger}_{i}\pm\hat{\tilde{c}}_{i}\right),$ (14)
We notice that the local direct and dual spaces are mixed in a simple way in each natural orbital, namely $\hat{c}^{\dagger}_{i}\pm\hat{\tilde{c}}_{i}$. We are allowed to perform a rotation of the super-orbitals such that the direct and dual spaces are not mixed anymore, one being associated to one natural orbital and one to the other:
$\hat{q}^{\dagger}_{\alpha,\mathrm{direct}}=\sum_{i=1}^{L}v_{i\alpha}^{*}\hat{c}^{\dagger}_{i}\quad;\quad\hat{q}^{\dagger}_{\alpha,\mathrm{dual}}=\sum_{i=1}^{L}v_{i\alpha}^{*}\hat{\tilde{c}}_{i}\,.$ (15)
The previous equation shows that the natural super-orbitals of a unitary operator are just a standard orbital rotation $v\hat{O}v^{\dagger}$. The super-orbitals $\hat{\hat{q}}$ are then simply the orbitals $\hat{q}$ associated to the space on which they act (direct or dual), such that the operator, which was vectorized as an MPS in a system with $2L$ orbitals, can again be written as an MPO in the super-orbitals basis. The following sketch illustrates the difference of the two types of rotations, in the super-orbital space for a generic operator (left panel) and in the standard orbital space for unitary operators (right panel).
Density matrices
We focus now on density matrices, especially for pure states as they are represented by rank one matrices. In this case, $R$ is completely equivalent to $Q$ for the corresponding states. The density matrix of a system in a pure state $|\psi\rangle$ can be written and vectorized:
$\hat{\rho}=|\psi\rangle\langle\psi|\rightarrow|\hat{\rho}\rangle\!\rangle=|\psi\rangle\otimes|\tilde{\psi}\rangle$. (16)
with the normalization $\langle\!\langle\hat{\rho}|\hat{\rho}\rangle\!\rangle=\langle\psi|\psi\rangle\langle\tilde{\psi}|\tilde{\psi}\rangle=1$. Note that for a pure state, $|\tilde{\psi}\rangle=|\psi\rangle$, but it does not matter to keep them different. The corresponding matrix $R$ reads, if we write it by blocks involving the original fermion operators as follows:
$R=\begin{pmatrix}\langle\psi|\hat{c}^{\dagger}_{m}\hat{c}_{n}|\psi\rangle&\langle\psi|\hat{c}^{\dagger}_{m}|\psi\rangle\langle\tilde{\psi}|\hat{\tilde{c}}_{n}|\tilde{\psi}\rangle\\ \langle\psi|\hat{c}_{n}|\psi\rangle\langle\tilde{\psi}|\hat{\tilde{c}}^{\dagger}_{m}|\tilde{\psi}\rangle&1-\langle\tilde{\psi}|\hat{\tilde{c}}^{\dagger}_{m}\hat{\tilde{c}}_{n}|\tilde{\psi}\rangle\end{pmatrix}$. (17)
The terms not preserving the particle number like $\langle\psi|\hat{c}^{\dagger}_{n}|\psi\rangle$ are simply 0, such that:
[…]
since $\tilde{Q}=Q$ for a pure state, the $R$ matrix is just twice the copy of the $Q$ matrix of the corresponding state when it’s particle-hole symmetric. For a mixed state holding a well defined charge, the matrices $Q$ and $\tilde{Q}$ are different but the dual and direct space do not mix as well. We can show the same results by using the definition of the elements of $R$ through the trace of operators as defined in Eq. (11) equivalently for density matrices. A pure state holds the following properties: $\hat{\rho}^{\dagger}=\hat{\rho}$ and $\hat{\rho}^{\dagger}\hat{\rho}=\hat{\rho}$. In this case, only the diagonal blocks of the $R$ matrix are non-zero contrary to unitary operators and they read:
$\mathrm{Tr}\left[\hat{\rho}\hat{c}^{\dagger}_{m}\hat{c}_{n}\right]=\langle\psi|\hat{c}^{\dagger}_{m}\hat{c}_{n}|\psi\rangle$ (19)
A simple example
In order to illustrate the notion of product operator, we provide here a simple example of an operator that can be factorized and whose natural super-orbitals occupations are either 0 or $1$. In general, an operator can be written as a sum of projectors between different product states in the basis it is defined, as we wrote in Eq. (5). Usually, operators are defined in the computational basis $|i\rangle$, in which each index corresponds to a fermion in a site: $|101\rangle=\hat{c}^{\dagger}_{2}\hat{c}^{\dagger}_{0}|0\rangle$. We take here the example of the operator $\hat{c}^{\dagger}_{0}$, which we aim to write in a factorized form, for $L=3$:
$\begin{split}\hat{c}^{\dagger}_{0}&=|001\rangle\langle 000|-|011\rangle\langle 010|\\ &-|101\rangle\langle 100|+|111\rangle\langle 110|.\end{split}$ (20)
We accounted for the fermionic signs, and the effect of the operator on an incoming state is to add a fermion at site 0 when allowed. Its vectorized form looks like:
$\begin{split}|\hat{c}^{\dagger}_{0}\rangle\!\rangle&=|001\tilde{0}\tilde{0}\tilde{0}\rangle\!\rangle-|011\tilde{0}\tilde{1}\tilde{0}\rangle\!\rangle\\ &-|101\tilde{1}\tilde{0}\tilde{0}\rangle\!\rangle+|111\tilde{1}\tilde{1}\tilde{0}\rangle\!\rangle\\ &=(\mathds{1}+\hat{c}^{\dagger}_{1}\hat{\tilde{c}}_{1}+\hat{c}^{\dagger}_{2}\hat{\tilde{c}}_{2}+\hat{c}^{\dagger}_{2}\hat{c}^{\dagger}_{1}\hat{\tilde{c}}_{2}\hat{\tilde{c}}_{1})\hat{c}^{\dagger}_{0}|0\rangle\!\rangle.\end{split}$ (21)
The vectorized operator can be normalized by adding a factor $1/2$. In the first line of the above expression, the dual and direct spaces are mixed and the tildes indicate the states from the dual space. The operator $\hat{c}^{\dagger}_{0}$ has a lot of redundancy, and it is possible to find a combination the creation and annihilation operators such that a more compact form can be written. We assume the following rotation:
$\begin{split}\hat{\hat{q}}^{\dagger}_{1}&=\hat{c}^{\dagger}_{0},\quad\hat{\hat{q}}^{\dagger}_{3}=\frac{-1}{\sqrt{2}}\left(\hat{c}^{\dagger}_{1}+\hat{\tilde{c}}_{1}\right),\quad\hat{\hat{q}}^{\dagger}_{5}=\frac{-1}{\sqrt{2}}\left(\hat{c}^{\dagger}_{2}+\hat{\tilde{c}}_{2}\right),\\ \hat{\hat{q}}^{\dagger}_{2}&=\hat{\tilde{c}}_{0},\quad\hat{\hat{q}}^{\dagger}_{4}=\frac{-1}{\sqrt{2}}\left(\hat{c}^{\dagger}_{1}-\hat{\tilde{c}}_{1}\right),\quad\hat{\hat{q}}^{\dagger}_{6}=\frac{-1}{\sqrt{2}}\left(\hat{c}^{\dagger}_{2}-\hat{\tilde{c}}_{2}\right).\end{split}$ (22)
In this basis, it is easy to reconstruct Eq. (21) from the new super-operators $\hat{\hat{q}}^{\dagger}_{m}$:
$|\hat{c}^{\dagger}_{0}\rangle\!\rangle=\hat{\hat{q}}^{\dagger}_{1}|0\rangle\!\rangle=|000001\rangle\!\rangle,$ (23)
In the new basis, the vectorized operator is now expressed in a single product state in the doubled Hilbert space. The rotation that we applied is unique (up to a phase), as it is the one diagonalizing the matrix $R[\hat{c}^{\dagger}_{0}]$. The spectrum of $R[\hat{c}^{\dagger}_{0}]$ in this case is then:
$\begin{split}\langle\!\langle\hat{\hat{n}}_{1}\rangle\!\rangle&=1,\langle\!\langle\hat{\hat{n}}_{3}\rangle\!\rangle=0,\langle\!\langle\hat{\hat{n}}_{5}\rangle\!\rangle=0,\\ 1-\langle\!\langle\hat{\hat{n}}_{2}\rangle\!\rangle&=1,\;1-\langle\!\langle\hat{\hat{n}}_{4}\rangle\!\rangle=1,\;1-\langle\!\langle\hat{\hat{n}}_{6}\rangle\!\rangle=1\\ &\equiv[0,0,1,1,1,1],\end{split}$ (24)
in which we can see that the occupations are $1$ away from half-filling, indicating that the operator $\hat{c}^{\dagger}_{0}$ creates one particle. Hence, in a similar sense as the matrix $Q$ is working for states, the power of the one-body density matrix $R$ is to reveal the basis in which a vectorized operator is written as a product state.
Note that the transformation above can be equivalently understood as by working in a Pauli-like basis $\{\mathds{1},\hat{c}_{i}^{\dagger},\hat{c}_{i},\hat{n}_{i}\}^{\otimes L}$ (mixing the local direct and dual spaces) for the operator, which reads $\hat{c}^{\dagger}_{0}\mathds{1}\mathds{1}$, being clearly a factorized form of the operator.
Correlation entropy
Similar to the case of states, we can define a single measure of the complexity of an operator based on the occupations of its natural super-orbitals. This measure is commonly referred to as the correlation entropy, and is defined analogously to the correlation entropy of states:
$S_{\mathrm{corr}}=-\sum_{\alpha=1}^{2L}n_{\alpha}\log\left(n_{\alpha}\right).$ (25)
When the operator can be factorized as a product operators, the occupations of the natural super-orbitals are either 0 or $1$, and the correlation entropy will be $S_{\mathrm{corr}}=0$. In the opposite case, the furthest from a product operator an operator is, the more natural super-orbitals have their occupation $n_{\alpha}$ close to $\frac{1}{2}$, and the correlation entropy will be maximal when every orbital reaches this value, leading to the extensive maximal correlation entropy $S_{\mathrm{max}}=L\log(2)$. We point out here that the spectrum of $R$ does not depend on the basis in which the operator is defined, operators which can be related through a one-body rotation (in the sense defined above) will have the same natural super-orbitals and thus the same correlation entropy.
Operator non-Gaussianity and non-stabilizerness
[…]
Non-stabilizerness: Characterizing the non-stabilizerness of an operator provides insight into its computational complexity, particularly in terms of tableau classical simulation [64, 38] and of the number of non-Clifford gates required for its realization on a quantum device. A standard measure of this resource is the Operator Stabilizer Rényi Entropy (OSE) [47, 46], defined as:
$\begin{split}\mathcal{M}^{(\alpha)}(\hat{O})&=\frac{1}{1-\alpha}\log\left(P^{(\alpha)}(\hat{O})\right)\\ P^{(\alpha)}(\hat{O})&=\sum\limits_{\mathcal{S}\in\mathcal{S}_{L}}\!\!\left(\frac{|\mathrm{Tr}\left(\hat{O}^{\dagger}\mathcal{S}\right)\!\!|}{2^{L}}\right)^{\!\!2\alpha}\!\!\!\!=\sum\limits_{\mathcal{S}\in\mathcal{S}_{L}}|a_{\mathcal{S}}|^{2\alpha},\end{split}$ (26)
where the coefficients $a_{\mathcal{S}}$ arise from the expansion $\hat{O}=\sum\limits_{\mathcal{S}\in\mathcal{S}_{N}}a_{\mathcal{S}}\hat{\mathcal{S}}$, which can be defined for the fermionic degrees of freedom as stated in the previous paragraph, $\mathcal{S}_{L}=\{\mathds{1},\hat{c}_{i}^{\dagger},\hat{c}_{i},\hat{n}_{i}\}^{\otimes L}$. Usually, $\alpha=2$ is considered. It measures the number of significant product operators (or Pauli strings when working in the computational basis) on which the operator is decomposed. In contrast to the non-Gaussianity, the non-stabilizerness is basis-dependent. For a Gaussian operator, we previously showed that it decomposes into a single product operator in the natural super-orbital basis, yielding vanishing non-stabilizerness. This basis also maximizes the weight of the coefficients $a_{\mathcal{S}}$ [59] given a number $r\leq 2L$ of super-orbitals included, thereby minimizing the non-stabilizerness under Gaussian transformations. Notably, when certain orbitals factorize in this basis, they do not contribute to the sum in Eq. (26), resulting in an exponential reduction in the number of terms and thus in the non-stabilizerness. Such behavior has been observed in the evolution of local operators that exhibit a light-cone structure, where orbitals outside the light-cone effectively decouple [58]. We will demonstrate that similar behavior arises in the time-evolution operator of impurity models when expressed in the natural super-orbital basis.
IV Special cases
We introduced above the theoretical framework for computing the one-body density matrix of an operator and illustrated its significance through a simple example. In this section, we apply the framework to two important cases frequently encountered in quantum many-body physics: the time-evolution operator of a non-interacting system and a Haar random unitary. These examples are particularly insightful because they reflect contrasting notions of complexity. In non-interacting systems, the Hamiltonian—and hence the time-evolution operator—can be diagonalized within the one-particle basis. In contrast, Haar random unitaries resemble infinite-temperature thermal states, where no simplification is expected from a one-body perspective. We demonstrate how both cases fit within the one-body density matrix framework: the time-evolution operator can be decomposed into a product of orbitals, while Haar unitaries exhibit maximal correlation entropy.
IV.1 Non-interacting systems
The Hamiltonian of a non-interacting system can be put in a diagonal form by diagonalizing its Hamiltonian matrix, which constitutes an orbital rotation. In the corresponding basis, it has the following local form:
$\hat{H}_{\mathrm{free}}=\sum_{n=1}^{L}\varepsilon_{n}\hat{n}_{n},$ (27)
where $\varepsilon_{n}$ and $\hat{n}_{n}$ are respectively the onsite energy and the occupation operator on the orbital $n$: each orbital is thus disconnected and all the terms commute with each other. The time-evolution operator reads:
$\hat{U}(t)=\exp\left(-it\sum_{n=1}^{L}\varepsilon_{n}\hat{n}_{n}\right)=\prod^{L}_{n=1}e^{-it\hat{n}_{n}},$ (28)
where the last equality uses the commutation relations of each term of $H$. We expand the exponential, and use the fact that $\hat{n}^{m}=\hat{n}$:
$\hat{U}(t)=\prod^{L}_{n=1}\left(\hat{\mathds{1}}+\hat{n}_{n}\left(e^{-it\varepsilon_{n}}-1\right)\right)=\prod^{L}_{n=1}\hat{O}_{n}.$ (29)
The form above is quite compact, and most importantly is a factorized form. Note that the vectorized form of the operator has to be normalized as previously stated. The vectorized form of $\hat{O}_{n}$ in its local basis reads:
$\begin{split}|\hat{O}_{n}\rangle\!\rangle&=|0\rangle\langle 0|+|1\rangle\langle 1|+(e^{-i\varepsilon_{n}t}-1)|1\rangle\langle 1|\\ &=(1+e^{-i\varepsilon_{n}t}\hat{c}^{\dagger}_{n}\hat{\tilde{c}}_{n})|0\rangle\!\rangle.\end{split}$ (30)
Similarly to Eq. (22), each term in the product can be rotated into the following pair of orbitals (since $\hat{U}(t)$ is unitary):
[…]
where $A=1-e^{-i\varepsilon_{n}t}$, such that $\hat{\hat{q}}^{\dagger}_{n}\hat{\hat{q}}^{\dagger}_{L+n}|0\rangle\!\rangle$ correctly reproduces Eq. (30). Note that these orbitals must be correctly normed. Each factorized operator $\hat{O}_{n}$ in Eq. (29) can thus be decomposed in two such kind of orbitals, which leads to:
$|\hat{U}(t)\rangle\!\rangle=\prod^{L}_{n=1}\hat{\hat{q}}^{\dagger}_{n}\hat{\hat{q}}^{\dagger}_{L+n}|0\rangle\!\rangle=|11\cdots 11\rangle\!\rangle.$ (32)
The spectrum of $R[\hat{U}(t)]$ consists of equal numbers of zeros and ones, as all orbitals are occupied —half of the spectrum corresponds to $\langle\!\langle\hat{\hat{n}}\rangle\!\rangle$, and the other half to $1-\langle\!\langle\hat{\hat{n}}\rangle\!\rangle$). Consequently, the operator $\hat{U}$ becomes trivial in a suitable super-orbital basis. Additionally, one can demonstrate that the operator $\hat{U}$ factorizes by directly evaluating matrix elements of $A$, defined in Eq. (13). Substituting Eq. (29) into this expression shows that $A_{mn}=\delta_{mn}\frac{e^{it\varepsilon_{n}}}{2}$. Using the relation between the spectra of $R$ and $A$ (i.e. $\lambda^{\pm}_{\alpha}=\frac{1}{2}\pm|\mu_{\alpha}|$), it follows immediately that $\lambda^{\pm}_{\alpha}\in\{0,1\}$.
The only condition imposed above is that the Hamiltonian must be quadratic in the fermionic operators. According to the Thouless theorem [65], any unitary operator representing a one-body rotation can be expressed as the exponential of a quadratic Hamiltonian. Hence, if the spectrum of the corresponding $R$ matrix consists only of zeros and ones, the associated unitary operator can be interpreted as an orbital rotation. However, this factorization does not extend to the Hamiltonian itself: as evident from Eq. (27), a sum of local operators cannot, in general, be written as a single product operator.
IV.2 Haar random operators
The case of random unitary operators is much simpler. It is possible to calculate the matrix elements $A_{mn}$ on average using the common machinery of Weingarten calculus [66], as shown in Appendix A. We find that for the time-evolution operator, these elements are all always zero, such that the spectrum of $R$ is fully degenerate with $\lambda=1/2$, corresponding to the worst case in which the maximal correlation entropy $S_{\mathrm{max}}=L\log(2)$ is reached, indicating maximal non-Gaussianity and non-stabilizerness. In such a case, there exist no one-particle basis in which a random unitary can be simplified in terms of any of the resources that we presented. As for $\hat{U}(t)$ the eigenvalues of $A$ are simply zero, we also investigate the behavior of the natural super-orbitals for a local operator which is rotated by a random unitary. In this case, we show that the eigenvalues of the matrix $A$ are decaying with the system size as $2^{-L}$ which converges exponentially to a maximally correlated operator.
V Results
We present the numerical study of the correlation spectrum for two specific unitary operators, the time-evolution operator $\hat{U}(t)$ and an evolved local operator of the type $(\hat{n}_{i}(t)-1/2)$ for two different models. The first model is the tVt-V model, in which every site is interacting with their neighbors, and the second one is the interacting resonant level model (IRLM) in which the interaction is localized at an edge of a tight-binding chain. Our numerical simulations are based on a $U(1)$-symmetric MPO representation of the corresponding operators, and employ the time-evolving block decimation (TEBD) algorithm [67, 9] in real space with a maximum bond dimension $\chi=1024$, as well as a $4^{\mathrm{th}}$-order Trotter decomposition with a time step $dt=10^{-2}$ to ensure the convergence of the data up to the times shown in the plots.
V.1 $t\text{-}V$ model
The first system investigated here is the tVt-V model of fermions in a one-dimensional open chain. This model is equivalent to the Heisenberg spin chain (or $XXZ$ model) through a Jordan-Wigner transformation [68], and is studied here as a prototypical interacting system, in which interaction and hopping compete at every site. The Hamiltonian is the following:
$\begin{split}\hat{H}_{\mathrm{t-V}}&=\sum_{i=1}^{L-1}\frac{\gamma}{2}\left(\hat{c}^{\dagger}_{i}\hat{c}_{i+1}+\hat{c}^{\dagger}_{i+1}\hat{c}_{i}\right)\\ &+V\left(\hat{n}_{i}-\frac{1}{2}\right)\left(\hat{n}_{i+1}-\frac{1}{2}\right),\end{split}$ (33)
in which the $1/2$ subtracted to density operators stems from the Heisenberg interaction $\hat{S}^{z}_{i}\hat{S}^{z}_{i+1}$ in the spin model, and enforces the particle-hole symmetry in the equation above. When studying the one-body correlation spectrum of operators associated with this model, it is important to emphasize that certain state-specific properties, which are typically well understood, do not always exist into this framework. Conventionally, the regimes $V>0$ and $V<0$ are associated with anti-ferromagnetic and ferromagnetic phases, respectively, in the underlying spin model, each exhibiting distinct physical behavior. However, within our formalism, these two regimes are entirely equivalent. Specifically, one can perform a basis rotation (as defined above) within the operator space $\{\mathds{1},\hat{c}^{\dagger}_{i},\hat{c}_{i},\hat{n}_{i}\}^{\otimes L}$, whereby all number operators $\hat{n}_{i}$ on even (or equivalently, odd) sites acquire a minus sign—e.g., $(\hat{n}_{0},\hat{n}_{1},\hat{n}_{2},\cdots)\rightarrow(-\hat{n}_{0},\hat{n}_{1},-\hat{n}_{2},\cdots)$. This transformation is formally equivalent to changing $V\rightarrow-V$. As discussed in Sec. III, the Hamiltonians $\hat{H}(V)$ and $\hat{H}(-V)$, as well as the corresponding time-evolution operator $\hat{U}(t)$, are related by such an orbital rotation and therefore share the same correlation spectrum. A similar argument applies to the sign of the hopping parameter $\gamma$, which can be reversed by applying a rotation to every other creation (or equivalently annihilation) operator $\hat{c}^{\dagger}_{i}$ (or $\hat{c}_{i}$).
Figure 1: Left: Spectrum of the one-body super-density matrix of the operators $\hat{U}(t)$ and $(\hat{n}_{25}(t)-1/2)$ for the $t$-$V$ model ($\equiv S^{z}_{25}(t)$ in the Heisenberg spin chain) with $L=50$ sites at different times. The local operator is taken in the middle of the chain at site $i=25$. The dashed line indicates $\lambda=0.5$, corresponding a maximally correlated orbital. Right: Correlation entropy of the corresponding operators. The parameters are indicated in the text.
In the left panel of Fig. 1, we show the spectra of $R[\hat{U}(t)]$ and $R[\hat{n}_{i}(t)-1/2]$, where $i$ is a site chosen at the center of a chain of length $L=50$, for time evolution with parameters $\gamma=V=1$. The blue curves correspond to the unitary operator $\hat{U}(t)$ at different times, initially represented as an identity MPO, to which Trotter gates are sequentially applied on one side to simulate time evolution. At $t=0$, the spectrum reflects a trivial operator, consisting solely of eigenvalues 0 and $1$. As time progresses, the spectrum evolves towards that of a highly correlated operator, with most orbital occupations approaching $1/2$. The spectral gap $\Delta$—defined as the difference between the largest occupation among those close to 0 and the smallest among those close to 11—initially equal to 1, closes over time in a non-algebraic manner. All orbitals become correlated similarly, lacking any notable internal structure except for the most and least occupied orbitals are affected by the open boundaries. The right panel of Fig. 1 presents the corresponding correlation entropy $S[\hat{U}(t)]$ as defined in Eq. (25), normalized by its maximal value $L\log(2)$. This entropy rapidly saturates, indicating that no preferential single-particle basis exists for $\hat{U}(t)$, and thus its structural complexity cannot be reduced via orbital rotations.
Figure 2: Spatial dispersion of the natural super-orbitals of an evolved local operator $\hat{n}_{L/2}(t)-1/2$ ($\equiv\hat{S}^{z}_{L/2}(t)$ in the spin model), for a chain of $L=50$ sites at time $t=8$. The overlap of the natural orbitals with real space orbitals is defined as the absolute value of the eigenvectors of the matrix $A$ as defined in Eq. (15), i.e. $|\langle\alpha|i\rangle|=|v_{i\alpha}|$. The full eigenspectrum of the matrix $R$ contains $2L$ orbitals, but the entire colormap consists simply four copies of the region displayed here.
The red curves illustrate the spectrum of $R[\hat{n}_{L/2}(t)-1/2]$ under the same conditions as for $\hat{U}(t)$. The initial operator is the identity at all sites except the center of the chain, where it is a local Pauli $S^{z}$ operator—thus also a trivial operator in terms of its natural super-orbitals. As time evolves, the spectrum deviates from the initial values 0 and $1$ in a structured manner: an increasing number of orbitals acquire occupations near $1/2$, while the rest remain close to 0 or $1$, with a smooth decay between the two regimes. This behavior reflects the light-cone structure characteristic of local operator spreading: at time $t$, the operator can be approximated as $\mathds{1}\otimes\cdots\otimes C\otimes\cdots\otimes\mathds{1}$, where $C$ denotes the non-trivial part of the operator within the light cone. If $C$ spans $c$ sites, then approximately $c$ orbitals exhibit occupations near $1/2$, while the remaining LcL-c remain close to 0 or $1$. The smooth decay between these regions arises because the operator is not an exact tensor product between $C$ and the outside region. This structure also governs the behavior of the correlation entropy, which grows approximately linearly in time at a significantly slower rate than that of the time evolution operator. It also bounds the growth of non-stabilizerness for such a local operator, which has been show to scale linearly in time [58] due to the factorized form, as explained earlier. At late times, when the support of $C$ extends across the entire system, the entropy eventually saturates near its maximal value.
A more detailed understanding of the structure of the correlation spectrum of $\hat{n}_{i}(t)-1/2$ can be obtained by examining the spatial dispersion of the natural super-orbitals. In Fig. 2, we display the absolute values of the coefficients $v_{i\alpha}$, defined in Eq. (15), which describe the spatial profile of the natural orbital $\hat{q}_{\alpha}$ at time $t=8$. The orbital indices $\alpha$ are ordered by increasing proximity of their occupation numbers to 0 or $1$, such that the most correlated orbitals—those with occupations closest to 0.5—appear first. At this time, the light cone emerging from site $L=25$ has propagated approximately $8$ sites in both directions, which is reflected in the complex, delocalized structure of the leading natural super-orbitals for $\alpha\lesssim 16$. As the occupation numbers approach 0 or $1$, the orbitals become increasingly localized, consistent with the exponential decay of correlations between regions inside and outside the light cone [69, 17]. Ultimately, each natural orbital becomes localized on a single site, in agreement with the cartoon of the operator structure discussed above, wherein sites outside the light cone are effectively in a tensor product state. Note that the maximum amplitude of each orbital is $1/\sqrt{2}$, as the diagonalization of the matrix $R$ yields natural super-orbitals defined in the doubled Hilbert space, as introduced in Eq. (14). Only the first $L$ orbitals (out of a total of $2L$) are shown in Fig. 2, since the remaining orbitals are redundant.
We illustrated here that bulk interacting models are not likely to be simplified through one-body orbital rotations. However, the correlation spectrum of local operators as well as the spatial dispersion of their natural super-orbitals exhibit footprints of underlying physical properties, which were until now investigated through different kind of mathematical objects such as OTOCs. In the following section, we demonstrate the power of the natural super-orbitals to reveal the hidden simplicity of some operators in specific models in which the interaction is localized on a few sites, namely quantum impurity models.
V.2 Impurity problem (IRLM)
The impurity model that we investigate in the following is the Interacting Resonant Level Model (IRLM), in which the spin-less nature of the fermions makes numerical simulations more practical. The model consists in two orbitals between which fermions can hop and interact (the impurity), coupled to a large tight-binding chain of non-interacting fermions (the bath). The Hamiltonian reads:
$\begin{split}\hat{H}_{\mathrm{IRLM}}&=U\left(\hat{n}_{1}-\frac{1}{2}\right)\left(\hat{n}_{2}-\frac{1}{2}\right)+V\left(\hat{c}^{\dagger}_{1}\hat{c}_{2}+\hat{c}^{\dagger}_{2}\hat{c}_{1}\right)\\ &+\frac{\gamma}{2}\sum_{i=2}^{L-1}\left(\hat{c}^{\dagger}_{i}\hat{c}_{i+1}+\hat{c}^{\dagger}_{i+1}\hat{c}_{i}\right),\end{split}$ (34)
where the $1/2$ subtracted to the density operators enforce particle-hole symmetry, and $\gamma/2$ is taken for the half-bandwidth of the bath to be $D=1$ for $\gamma=1$. Hence we work in units of $D$ (or equivalently of $\gamma$) in the following. As for the tVt-V model presented above, the correlation spectrum is symmetric under super-orbital rotations, which in this model also corresponds to flipping the signs of the parameters: $U\rightarrow-U$, as well as $V\rightarrow-V$ or $\gamma\rightarrow-\gamma$ yields the same properties of the correlation spectrum. In the IRLM, the ground state properties of the model with $U>0$ and $U<0$ are very different, as the Kondo temperature is exponentially smaller for negative values of the interaction, leading to a more correlated state [30]. Such an effect is not present here as the properties of the operator are related to its full spectrum, such that specific properties related to exceptional states (like the ground state) are completely diluted. In the following, we first discuss the properties of the time-evolution operator $\hat{U}(t)$ for different interaction strengths $0<U<D$ and $U>D$, and then we study the properties of an evolved local operator on the impurity site $\hat{n}_{1}(t)-1/2\equiv\hat{S}^{z}_{1}(t)$. The numerical parameters that we choose for our simulations are the same as the ones used for the Heisenberg model.
Figure 3: Left: Correlation spectrum of the time evolution operator $\hat{U}(t)$ at different times for the IRLM, showing the half of orbitals the closest to 0, the other half being the particle-hole symmetric $1-\lambda_{\alpha}$. The plateau arising after $t\geq 8$ is due to the truncation error, as the corresponding MPO has reached the maximum bond dimension $\chi=1024$. The $x$-axis is cut at $\alpha=16$ instead of $\alpha=2L$ for visibility, as the remaining eigenvalues are below the machine precision. Right: Correlation entropy of $\hat{U}(t)$ as a function of time for different values of the interaction in the regime $U<D$, normalized by the maximal extensive value $S_{\mathrm{max}}=L\log{2}$. For both panels, we consider as system of $L=50$ sites. For both panels, $V=0.2D$.
Time-evolution operator: The correlation spectrum of the operator $\hat{U}(t)$ is shown in the left panel of Fig. 3 for $U=0.3D$ in a system with $L=50$ sites. We present the spectrum at various times up to $t=9$, beyond which a plateau appears in the spectrum around $10^{-6}$, attributable to truncation errors. The figure displays only the part of the spectrum near 0, as the full spectrum is symmetric about $1$. The key observation is that, similar to the ground state of this model [30, 31], the occupations of the natural super-orbitals of the time-evolution operator decay exponentially at all times. However, this decay becomes progressively slower as time increases, eventually leading the spectrum to converge toward the maximal value of the correlation entropy $S_{\mathrm{max}}=L\log(2)$.
Figure 4: Correlation entropy of the time-evolution operator $\hat{U}(t)$ against time for the IRLM with $L=50$ sites, in the large-interaction regime $U>D$. The correlation entropy is normalized by its maximal value $S_{\mathrm{max}}=L\log{2}$. The gray curve labeled by $U\to\infty$ corresponds to the analytical result at first order in $1/U$, for $U=20D$.
The right panel of Fig. 3 shows the time evolution of the correlation entropy for different interaction strengths in the weak-coupling regime $U<D$. After a timescale comparable to the electronic bandwidth $D$, the correlation entropy begins to grow linearly with time. As the interaction strength increases, the entropy grows more rapidly, indicating that the operator becomes more entangled—i.e., the occupations of a larger number of natural super-orbitals deviate more significantly from 0 and $1$. Nonetheless, the absolute value of the correlation entropy remains well below the maximal limit, implying that the operator admits a relatively compact representation in the natural orbital basis. In contrast, its real space MPO representation used in these simulations reaches the imposed bond dimension $\chi=1024$ already at $t=8$, due to the much faster growth of the von Neumann entropy in that basis. Rotation to natural super-orbitals thus offers great compression of the considered evolution operators.
The linear growth of $S_{\mathrm{corr}}$ indicates that the non-Gaussianity of $\hat{U}(t)$ increases steadily over time. Additionally, the correlation spectrum allows us to estimate an upper bound on the non-stabilizerness generated by time evolution under a quantum impurity Hamiltonian. At any given time, the occupations of the natural super-orbitals decay exponentially. Therefore, orbitals with occupations below a given threshold $\delta$ can be neglected, as they effectively factorize and do not contribute to the measure of non-stabilizerness. We observe that the number of orbitals with occupations $n_{\alpha}>\delta$ increases linearly with time as $2vt$, where the factor $2$ comes from the almost occupied and empty orbitals coming together and $v$ is a parameter dependent factor, which is larger for larger interaction $U$. Consequently, the number of Pauli strings with non-negligible coefficients (for arbitrarily small $\delta$) in the natural super-orbital basis grows proportionally to $c2^{2vt}$, where $c$ is a constant depending on system parameters and $\delta$. In the worst-case scenario, where all Pauli strings contribute equally with $a_{S}=1$, the non-stabilizerness of the time-evolved operator is bounded by:
$\mathcal{M}_{\mathrm{NSO}}(\hat{U}(t))\leq 2vt\log(2)+\tilde{c}$, (35)
where $\tilde{c}$ is a constant and the subscript NSO refers to the natural super-orbital basis. This bound can be further tightened by accounting for the exponential decay of super-orbital occupations above the threshold $\delta$, which imposes additional structure on the coefficients $a_{S}$. Equation (35) resembles the known bound on the growth of magic for local operators, which also scales linearly in time [58] due to the factorization of real-space orbitals outside the light cone. It is somewhat surprising that the time-evolution operator $\hat{U}(t)$ displays a light-cone-like structure, given that it is completely non-local in real space. However, in the natural super-orbital basis, the majority of orbitals factorize due to the locality of the underlying interaction and thus do not contribute significantly to the Pauli decomposition. As time evolves, local interactions cause correlations to spread across more orbitals, increasing the number of non-negligible Pauli strings in a manner reminiscent of a light-cone structure, despite the operator’s inherent non-locality.
The time evolution of the correlation entropy in the large-interaction regime $U>D$ is shown in Fig. 4. For moderate interaction strength $U=2D$, close to the bandwidth $D$, the correlation entropy continues to increase with time, but superimposed oscillations begin to emerge. As $U$ increases further, these oscillations become more pronounced, with higher frequencies and amplitudes bounded by $2\log(2)$. This behavior can be understood by considering the limit $U\to\infty$, where the impurity becomes effectively decoupled from the bath. In this limit, the interaction at the impurity no longer influences the bath orbitals, which consequently factorize and do not contribute to the correlation entropy. The relevant degrees of freedom are then restricted to the two impurity sites, giving rise to four non-trivial natural super-orbitals. In the decoupled limit, the occupations of these natural super-orbitals exhibit oscillatory dynamics with distinct frequencies, governed by expressions of the form:
$\langle\!\langle\hat{n}_{i}(t)\rangle\!\rangle_{U\to\infty}=\frac{1}{2}\pm\frac{1}{2}\cos\left((V\pm\frac{U}{2})\,t\right),$ (36)
see Appendix B for details on the derivation. These slightly detuned frequencies produce a correlation entropy that oscillates with two characteristic frequencies: a fast component $\omega_{1}=2/U$, and a slow component $\omega_{2}=1/V$. Additionally, this toy model indicates that spatial profiles of the natural super-orbitals oscillate between the two impurity sites. When one site is fully occupied and the other is empty, the orbitals factorize and the correlation entropy vanishes. Conversely, when both sites are half-filled, the orbitals are maximally entangled, resulting in the maximum entropy for four orbitals, $2\log(2)$. The analytical expression for the correlation entropy at first order in $1/U$ corresponds to the gray curve in Fig. 4, and the rescaled maximal value $2/L$ is indicated by a dashed line. The fast oscillations are correctly captured by the frequency $\omega_{1}$ for $U=20D$. This is not the case for the slow component: in Fig 4, using the simulation’s parameter $V$ leads to a faster oscillating envelope, which we had to rescale here to match the data. In the toy model, the impurity is completely decoupled to the bath and is then agnostic to any bandwidth, which must play a role in the slow oscillating component related to low energies.
Figure 5: Left: Occupation of the natural super-orbitals of $\hat{n}_{1}(t)-1/2$ in the IRLM for $L=100$ and $V=U=0.2D$ at different times, plotted as in Fig. 3. Right: Correlation entropy of the same operator against time for different interaction strengths. For larger interaction strengths (keeping $U<D$), the correlation entropy saturates later in time at a larger value.
Local operator: We now investigate the properties of the time-evolved local operator, focusing on the $U<D$ regime, where impurity-bath correlations are not spoiled by local effects. The correlation spectrum of $\hat{n}_{1}(t)-1/2$ is shown in the left panel of Fig. 5 as a function of time, for system size $L=100$, and parameters $V=U=0.2D$, up to time $t=50$. The light-cone structure simplifies the MPO representation of this operator, such that no truncation effects are observed for these parameters. Nevertheless, the bond dimension keeps increasing and reaching later times is not possible.
Our first observation is that the natural orbital occupations of this operator decay exponentially as for $\hat{U}(t)$. However, the structure of the correlation spectrum is richer: we identify two distinct groups of natural super-orbitals—one with relatively large occupations, and a second with significantly smaller occupations with an elbow-like feature. As time progresses, orbitals from the second group transition into the first, populating the tail of the exponential decay. This behavior is illustrated in more detail in Fig. 6, where the first and second groups are highlighted by red and orange rectangles, respectively.
Figure 6 presents the time evolution of the occupations of every other natural super-orbitals, clearly showing the gradual transfer from the second group to the first. The increase in occupation of these orbitals reflects the expansion of the light-cone structure of the operator: as each new site becomes involved, it contributes a correlated orbital to the first group. Each newly included orbital exhibits an occupation that is exponentially smaller than the previous ones, with the decay rate governed by the interaction strength $U$. As with the time-evolution operator, a larger $U$ results in a slower decay, and thus a more correlated operator.
Figure 6: Time evolution of the occupations of the natural super-orbitals of the local operator $\hat{n}_{1}(t)-1/2$ for $L=100$, $V=0.2D$ and $U=0.1D$. A smaller interaction strength $U$ is chosen to appreciate a faster exponential decay of the largest eigenvalues. The colored rectangles indicates the two groups of natural super-orbitals based on their occupation.
Eventually, the gap between the two groups vanishes, signaling a saturation in the complexity of the operator. This saturation is captured in the right panel of Fig. 5, which shows the evolution of the correlation entropy. The point of saturation appears to depend on the initial gap between the two orbital groups—related to the ratio U/VU/V—and the rate of exponential decay within the first group, which is controlled by $U/D$.
The spatial distribution of the natural super-orbitals associated to the local operator $\hat{n}_{1}(t)-1/2$, as depicted in Fig. 7, reflects the hierarchical organization of orbital correlations, which decay exponentially. As previously discussed in the context of Fig. 2, the most strongly correlated orbitals correspond to indices $\alpha\lesssim 30$. In contrast, for large values of $i$ and $\alpha$, the eigenvectors of the matrix $R$ appear effectively random. This is a consequence of the associated eigenvalues falling below machine precision, rendering them numerically degenerate. These orbitals lie outside the light-cone and, when projected onto the real-space basis, resemble the structure observed in the exterior region of the light-cone in Fig. 2.
The two aforementioned distinct groups of orbitals can also be identified by their spatial dispersion. The first group, composed of the most strongly correlated orbitals, exhibits a delocalized structure. As the orbital correlation strength decreases (corresponding to increasing $\alpha$), these orbitals progressively localize, eventually leading to an orbital predominantly localized at the first site. Beyond this point, the second group of orbitals—characterized by much lower occupations emerges, displaying localization patterns around sites increasingly distant from the impurity following a linear light-cone-like propagation. This behavior suggests distinct roles for the two orbital groups. The highly correlated, delocalized orbitals encode the strong, shorter-ranged correlations induced by the local impurity interaction, akin to the formation of a Kondo cloud in the ground state of quantum impurity models. In contrast, the second group captures the propagation of the local perturbation through the system via nearest-neighbor hopping, representing the light-cone structure of the time-evolved operator. At long times and for sufficiently large systems, these orbitals become effectively localized on individual sites, indicating that their dynamics are no longer influenced by the impurity. This supports the observation that the correlation entropy of the local operator saturates in time, even as system size increases. The spatial extent of the most correlated orbital increases with time until it saturates at a finite length scale, which itself grows with the interaction strength $U$ (in the regime $U<D$). Both the spatial spread and the number of orbitals in this highly correlated group increase with $U$ in this regime. The other regime $U>D$ is less interesting, as the most correlated orbitals become more localized near the impurity, and the second group of orbitals becomes dominant.
Figure 7: Spatial dispersion of the natural super-orbitals $q_{\alpha}$ of the evolve density operator on the impurity $\hat{n}_{1}(t)-1/2$, for a chain of $L=100$ sites at time $t=30$, for $U=V=0.2D$. Only the first $\approx 30$ eigenvalues are above machine precision, such that the eigenvectors of the rest are random due to spurious degeneracies.
V.3 Early growth of the correlation entropy
Finally, we study the short time behaviour of the operators studied above for the two models presented here, the tVt-V (or Heisenberg) model and the IRLM. At short times, the elements of the matrix $R$ of these unitary operators defined in Eq. (11) can be approximated using the following form of the Baker-Hausdorff-Campell (BHC) formula:
$e^{\hat{X}}\hat{Y}e^{-\hat{X}}=\sum_{n=0}^{\infty}\frac{1}{n!}[\hat{X},\hat{Y}]_{p},$ (37)
where $\hat{X},\hat{Y}]_{p}=[\hat{X},[\hat{X},\hat{Y}]]_{p-1}$ and $\hat{X},\hat{Y}]_{0}=\hat{Y}$.
We first consider the time-evolution operator $\hat{U}(t)=e^{-i\hat{H}t}$ and compute the matrix elements of $A$ to leading orders in $t$. In this framework, we define $\hat{X}=i\hat{H}t$ and $\hat{Y}=\hat{c}_{m}^{\dagger}$, while $\hat{c}_{n}$ remains time-independent and serves as a spatial probe for the spread of $\hat{c}_{m}^{\dagger}$ under time evolution, in a manner analogous to an OTOC.
Figure 8: Correlation entropy of the time-evolution operator $\hat{U}(t)$ (upper panel) and of an evolving local operator $\hat{n}_{i}-1/2$ (lower panel), for chains of $L=50$ in the tVt-V model (blue curves) and in the IRLM model (red curves). The dashed curves indicate the expected scalings from perturbation theory (see text).
For the Heisenberg model, the diagonal elements of $A$ are constant at leading order, with $A_{mm}=1/2$. For the off-diagonal elements, we find that $\hat{U}^{\dagger}\hat{c}_{m}^{\dagger}\hat{U}$ commutes with $\hat{c}_{n}$ up to order p=|nm|p=|n-m| in the BCH expansion, implying $A_{mn}\propto t^{|n-m|}$. Consequently, retaining only the leading order contribution for |nm|=1|n-m|=1 reduces $A$ to a tridiagonal form. Following Eq. (13), we compute the spectrum of $A^{\dagger}A$ which is also a tridiagonal matrix. Its spectrum can then obtained analytically and is given by $\mu_{\alpha}=1/2-t^{2}c_{\alpha}/2$, where the coefficients $c_{\alpha}$ encode spectral details independent of $t$. Accordingly, the spectrum of $R[\hat{U}(t)]$ is given by $\lambda_{\alpha,+}=1-t^{2}c_{\alpha}$ and $\lambda_{\alpha,-}=t^{2}c_{\alpha}$. Due to the isotropy of the Heisenberg model, all eigenvalues exhibit the same scaling behavior, leading to an extensive number of natural super-orbitals becoming more correlated at the same time.
In contrast, for the IRLM, the interaction anisotropy causes the diagonal elements of $A$ in the bath to vanish as $t^{|m|}$ for $m>2$, so that only the impurity contributions remain dominant at leading order. In this case, the scaling of eigenvalues remains $\lambda_{\alpha,+}=1-t^{2}c_{\alpha}$ and $\lambda_{\alpha,-}=t^{2}c_{\alpha}$, but applies only to the leading modes ($\alpha=1,2$), while all other eigenvalues are either 0 or $1$. These behaviors are illustrated in the upper panel of Fig. 8, which shows the correlation entropy for both models across different parameters for a system size of $L=50$. As predicted, the correlation entropy grows quadratically in time ($\propto t^{2}$) for both models. However, the amplitude is larger in the Heisenberg case, reflecting the fact that more natural super-orbitals contribute at the same leading order.
The case of the local operator $\hat{n}_{i}(t)-1/2$ requires a slightly different treatment. As before, we set $\hat{X}=i\hat{H}t$, but now choose $\hat{Y}=\hat{n}_{i}$. For the tVt-V (or Heisenberg) model, the site index $i$ is taken at the center of the chain ($i=L/2$), while for the IRLM, $i=1$ corresponds to the impurity site. In both models, the diagonal and off-diagonal elements of the correlation matrix decay faster than $1/t$ as their distance from site $i$ increases. Consequently, the dominant contributions at leading order arise from a reduced $(3\times 3)$ sub-matrix of $A$, corresponding to indices m,n=i1m,n=i-1 to $i+1$. The structural differences between the two Hamiltonians lead to distinct scaling behaviors of the spectrum of $A$: in the Heisenberg model, the spectrum scales as $S_{\mathrm{corr}}\propto t^{4}$ (first non-vanishing terms), whereas in the IRLM it scales as $S_{\mathrm{corr}}\propto t^{6}$. These predictions are confirmed by numerical simulations, as shown in the bottom panel of Fig. 8.
This highlights that the short-time scaling of the correlation spectrum of the time-evolution operator may be insensitive to fine details of the Hamiltonian. In contrast, evolving local operators provide a more refined probe, capturing model-specific features. This sensitivity is reminiscent of the behavior of OTOCs, which also involve the dynamics of local operators.
VI Conclusion and outlook
In this work, we introduced the concept of natural super-orbitals associated with many-body operators, defined as the eigenvectors of the one-body super-density matrix constructed from a vectorized operator. To illustrate the meaning of these orbitals in the super-Hilbert space of $2L$ particles, we analyzed several simple cases. We demonstrated that a suitable super-orbital rotation can render an operator trivial in the MPO representation for non-interacting systems, while no such simplification is possible for random operators. We introduced measures of non-stabilizerness and non-Gaussianity for operators and discuss their properties related to natural super-orbitals. Additionally, we showed that for unitary operators, a super-orbital rotation corresponds to a standard basis change that preserves the meaning of the operator in the physical Hilbert space of $L$ particles.
We then applied this framework to characterize the natural super-orbitals of both the time-evolution operator and a time-evolved local operator in two interacting models: the tVt-V model and a quantum impurity model. In the impurity model, the natural super-orbitals reveal a compressed structure in both operators, evident through the exponential decay of their occupation numbers. This suggests that expressing these operators in the natural orbital basis allows for a significantly more compact MPO representation at all times, enabling the study of larger systems and longer time scales beyond the reach of current real-space tensor network methods. We also discussed the implication of the exponential decay in terms of quantum resources, and showed that the non-stabilizerness of the time evolution operator of the impurity model has light-cone-like properties in the natural super-orbital basis stemming from the locality of the interaction. Using the correlation entropy, we observed that the complexity or non-Gaussianity of a local operator initially located on the impurity site saturates over time, with the saturation timescale increasing with interaction strength. This behavior suggests that quantum correlations becomes absent at large scales, even at late times.
Future directions include the development of a few-body detection algorithm, analogous to that introduced for quantum states in Ref. [24], to efficiently simulate OTOCs in quantum impurity models at long times and for large systems, in order to test the Lieb-Robinson bound in these systems. Such a tool would validate and extend the present findings, shedding light on the scrambling behavior in these systems. Finally, this work opens up new perspectives for studying orbital correlations in operator space across different physical contexts—for example, in many-body localized systems, where natural orbitals associated with quantum states have shown intriguing properties [32, 33, 34, 35]. Another promising avenue is the investigation of natural super-orbitals in mixed-state density matrices represented as MPOs, aiming to understand the impact of interaction with the environment and to provide a one-body framework for analyzing the loss of correlations due to dissipation.
Acknowledgement. The author thank Serge Florens, Yuriel Núñez-Fernández, Xhek Turkeshi and Frank Pollmann for interesting discussions. The author was supported by the Deutsche Forschungsgemeinschaft through the cluster of excellence ML4Q (EXC 2004, project-id 390534769) and by the Deutsche Forschungs- gemeinschaft through CRC 1639 NuMeriQS (project-id 511713970).
Appendix A Weingarten calculation for Haar random unitaries
In this appendix, we detail the derivation of the correlation spectrum for Haar random unitaries. We consider the elements of the matrix $A$ as the averaged value over the Haar measure of the $d$-dimensional unitary group $U(d)$, first for the chosen operator to be the time-evolution operator:
$\mathbb{E}_{U}(A_{mn}[\hat{U}])=\int\mathrm{d}U\,\mathrm{Tr}\left[U^{\dagger}\hat{c}^{\dagger}_{m}U\hat{c}_{n}\right],$ (38)
and also for an evolving local operator $\hat{O}=U^{\dagger}\hat{n}_{i}U$:
$\mathbb{E}_{U}(A_{mn}[\hat{n}_{i}])=\int\mathrm{d}U\,\mathrm{Tr}\left[U^{\dagger}\hat{n}_{i}U\hat{c}^{\dagger}_{m}U^{\dagger}\hat{n}_{i}U\hat{c}_{n}\right],$ (39)
The average over the unitary group is performed using Weingarten calculus, using its graphical representation. Following Ref. [66], we can express the expected value in terms of diagrams $\mathcal{D}_{r}$ obtained after a removal procedure $r$ of the original diagram $\mathcal{D}$, chosen in the space of possible removals $r\in\mathrm{Rem}(\mathcal{D})$. The expectation value thus reads:
$\mathbb{E}_{U}(\mathcal{D})=\sum_{r=(\alpha,\beta)\in\mathrm{Rem}(\mathcal{D})}\mathcal{D}_{r}\mathrm{Wg}(d,\alpha\beta^{-1}),\$(40)
where $\mathrm{Wg}(d,\sigma)$ is the Weingarten function, $d$ is the dimension of the matrix $U$, and $(\alpha,\beta)$ is a set of two permutations defining the corresponding removal $r$, which we explain further below. In the large-$d$ limit, the Weingarten function takes simple forms that we directly plug in our equations [66, 70].
Figure 9: Weingarten diagram for the trace of operators $\mathrm{Tr}\left(\hat{O}\hat{c}^{\dagger}_{m}\hat{O}\hat{c}_{n}\right)$, corresponding to the elements of the matrix $A$ for Haar random unitaries. a) $\hat{O}=\hat{U}(t)$. b) $\hat{O}=\hat{U}^{\dagger}(t)(\hat{n}_{k}-1/2)\hat{U}(t)$. We note $U^{\dagger}=\bar{U}^{t}$, where the transpose is performed by inverting the direct and dual spaces, denoted by empty and full circles respectively. Blue lines connect direct spaces of $U_{i}$ and $U_{\alpha(i)}$ and red lines connect the dual spaces of $U_{i}$ and $U_{\beta(i)}$. Solid and dashed lines correspond to the two different possible permutations for $\alpha$ and $\beta$.
Removals: In order to explain the way to perform the removal, we refer to Fig. 9, which shows a diagrammatic representation of Eq. (38) and Eq. (39) (panels a) and b) respectively). In these diagrams, white (resp. black) circles refer to the direct (dual) Hilbert space on which an operator (represented by boxes) act. We also performed the transpose of $\bar{U}=(U^{\dagger})^{t}$ by inverting black and white circles on the corresponding operators. In this case, $U_{1,2}$ and $\bar{U}_{1,2}$ correspond to a same random realization of a random unitary matrix $U$, but the indices are useful to define a removal $r=(\alpha,\beta)$. $\alpha$ (resp. $\beta$) is a permutation connecting the white (black) circle of $U_{i}$ to the white (black) circle of $\bar{U}_{\alpha(i)}$ ($\bar{U}_{\beta(i)}$). Each combination of $\alpha$ and $\beta$ corresponds to a removal $r$: we connect the lines (vector spaces) according to the chosen permutations, and read what is left in $\mathcal{D}_{r}$.
In the panel a) of Fig. 9, there is only a single choice for $\alpha$ and $\beta$, leading to a single reduced diagram. The reduced diagram reads $\mathcal{D}_{r}=\mathrm{Tr}(\hat{c}^{\dagger}_{m})\mathrm{Tr}(\hat{c}_{n})=0$, since these operator only have off-diagonal elements. The elements of $A$ are thus always 0, such that the eigenvalues of $R$ are trivially all equal to $1/2$.
In the panel b) of Fig. 9, we have two choices for $\alpha$ (blue lines) and two for $\beta$ (red lines), which leads to four reduced diagrams. We summarize the different combinations in the following table:
$\begin{matrix}\alpha&&\beta&&\mathcal{D}_{r}&&\mathrm{Wg}(d,\alpha\beta^{-1})\\ \hline\cr(12)&&(1)(2)&&\mathrm{Tr}(\hat{c}^{\dagger}_{m})\mathrm{Tr}(\hat{c}_{n})\mathrm{Tr}(\hat{O})^{2}&&-d^{-3}\\ (12)&&(12)&&\mathrm{Tr}(\hat{c}^{\dagger}_{m})\mathrm{Tr}(\hat{c}_{n})\mathrm{Tr}(\hat{O}^{2})&&d^{-2}\\ (1)(2)&&(1)(2)&&\mathrm{Tr}(\hat{c}^{\dagger}_{m}\hat{c}_{n})\mathrm{Tr}(\hat{O})^{2}&&d^{-2}\\ (1)(2)&&(12)&&\mathrm{Tr}(\hat{c}^{\dagger}_{m}\hat{c}_{n})\mathrm{Tr}(\hat{O}^{2})&&-d^{-3}\end{matrix},$
where $\hat{O}=\hat{n}_{i}-1/2$. Only $\mathrm{Tr}(\hat{c}^{\dagger}_{m}\hat{c}_{n})$ and $\mathrm{Tr}(\hat{O}^{2})$ are non zero, such that only the last element of the table above is not zero. This last term evaluates to $\mathrm{Tr}(\hat{c}^{\dagger}_{m}\hat{c}_{n})\mathrm{Tr}(\hat{O}^{2})\mathrm{Wg}(d,\alpha\beta^{-1})=-\delta_{mn}\frac{d}{2}\frac{d}{4}d^{-3}=-\delta_{mn}\frac{1}{8d}$. Hence the eigenvalues of $R$ are given by $\lambda_{n}=1/2\pm 1/(8d)$, where $d=2^{L}$ is the dimension of the corresponding Fock space. Local operators rotated by a random Haar unitary have therefore a non trivial correlation spectrum for all finite system sizes, but tend exponentially to the one of a fully correlated operator with system size.
Appendix B Correlation spectrum in the atomic limit
In the limit of large interaction strength $U\to\infty$, we assume that the impurity is disconnected from the bath and study the correlation matrix in the atomic limit. The Hamiltonian that we consider is:
$H=U\left(\hat{n}_{1}-\frac{1}{2}\right)\left(\hat{n}_{2}-\frac{1}{2}\right)+V\left(\hat{c}^{\dagger}_{1}\hat{c}_{2}+\hat{c}^{\dagger}_{2}\hat{c}_{1}\right),$ (41)
and the corresponding time-evolution operator takes the form of a $(4\times 4)$ matrix:
[…]
where $\omega=V/U$ and $\tilde{t}=Ut$. We can compute the elements of the matrix $A$ defined as $A]_{mn}=\mathrm{Tr}[\hat{U}^{\dagger}(t)\hat{c}^{\dagger}_{m}\hat{U}(t)\hat{c}_{n}]/\mathrm{Tr}[\mathds{1}]$, which reads:
$A=\frac{1}{2}\begin{bmatrix}\cos(\omega\tilde{t})\cos(\frac{\tilde{t}}{2})&-\sin(\omega\tilde{t})\sin(\frac{\tilde{t}}{2})\\ -\sin(\omega\tilde{t})\sin(\frac{\tilde{t}}{2})&\cos(\omega\tilde{t})\cos(\frac{\tilde{t}}{2})\end{bmatrix}$. (43)
It is straightforward to find the eigenvalues of the matrix $A^{\dagger}A$, which are the following oscillating functions:
$\mu_{\pm}=\cos(\omega\tilde{t}\pm\frac{\tilde{t}}{2})^{2},$ (44)
from which we find the four eigenvalues of $R$ as $\lambda_{n}=1/2\pm\sqrt{\mu_{n}}=1/2\pm\cos((V\pm U/2)t)$, as presented in Eq. (36). Note that this equation allows to find the correct fast component of the oscillations, which is given by the dominating term $U$. However, the slow component is not correctly captured as it would require to introduce effects from the bath which are of the same order in $1/U$, since $V$ is not considered large here.